# Bias in Large Language Models: Origin, Evaluation, and Mitigation

### v1

@@ 4 Bias Evaluation
As LLMs become increasingly integrated into various real-world applications, from healthcare decision-making to legal judgments and everyday digital interactions, their potential to propagate biases has raised significant ethical and societal concerns Liang et al. (2021). These biases, deeply rooted in the data, model architecture, and even in post-processing steps, can lead to discriminatory outcomes against marginalized groups, amplifying existing inequalities Gallegos et al. (2024). Therefore, understanding and evaluating these biases are crucial steps in ensuring the fair and responsible deployment of LLMs.
@@ 4.1 Data-Level Bias Evaluation Methods
Data-level bias evaluation methods focus on identifying and quantifying biases inherent in the training data of LLMs. These biases, rooted in the data, can significantly impact model outputs and exacerbate societal stereotypes.
@@
    • Representation Across Demographics: Research has consistently demonstrated that the representation of demographics in training data can significantly influence the outcomes of LLMs. For instance, LLMs trained on datasets with disproportionate representation of certain demographic groups are prone to exhibit biases that favor those groups Simmons and Hare (2023); Wang et al. (2024a); Gorti et al. (2024). For example, an analysis of political biases in LLMs revealed that most conversational LLMs exhibit left-of-center political preferences when probed with politically charged questions or statements Rozado (2024). The study demonstrated that LLMs can be steered towards specific political orientations through Supervised Fine-Tuning (SFT) with modest amounts of politically aligned data.
      To assess representation, statistical tools can be employed to measure the distribution of demographic attributes within the dataset. Key metrics might include the relative frequency of each demographic group or the variance in representation across different categories. Visual tools such as histograms, bar charts, or demographic distribution tables can be useful for identifying disparities, providing a clear picture of which groups are overrepresented or underrepresented in the data.
@@
    • Data Sources Analysis: The quality and bias of a model are heavily influenced by the sources of the data used for training. Biases inherent in the data sources can significantly impact the model’s behavior. For instance, if a model is trained primarily on data from Western countries, it may perform poorly when applied in non-Western contexts.
@@
      Evaluating the bias of data sources for LLM training involves analyzing the origins, diversity, and quality of the data, as well as its impact on model performance. This includes cataloging the data sources to ensure they cover a wide range of geographical, cultural, and linguistic contexts, and examining the credibility and inherent biases of each source. For instance, research Shen et al. (2023) on the SlimPajama dataset, which includes a rigorously deduplicated combination of web text, Wikipedia, GitHub, and books, reveals how different data combinations affect LLM performance. The SlimPajama-DC study highlights two key aspects: the impact of global versus local deduplication on model performance and the importance of data diversity post-deduplication. Findings indicate that models trained with highly-deduplicated, diverse datasets outperform those trained on less refined data, underscoring the significance of comprehensive and balanced data sources. Bias detection tools and adjustments in data weighting can further help in managing these biases, supported by transparent documentation and regular reviews to ensure fairness and accuracy in the model outputs.
@@
  • Stereotype and Bias Detection in Text Data: Stereotype and bias detection involves analyzing the content of the training data to identify and quantify biases related to stereotypes, offensive language, or prejudiced statements. This can be done through:
@@
    • Lexical Analysis: This method focuses on identifying specific words or phrases within the training data that are associated with bias or stereotypes. Lexical analysis relies on pre-defined lexicons, which are curated lists of terms known to carry biased or stereotypical connotations. Tools like Hatebase, an extensive repository of hate speech terms, provide a valuable resource for identifying harmful language in text data. Hate speech can also be measured by large scale study Silva et al. (2016). BiCapsHate Kamal et al. (2023) is deep learning model to detect hate speech in online social media posts. Similarly, the Dictionary of Affect in Language (DAL) Whissell (1989) categorizes words based on their emotional connotations, offering insights into how language can perpetuate stereotypes. By scanning the training data for occurrences of these terms, researchers can quantify the prevalence of biased language and identify specific areas where the data may reinforce harmful stereotypes. For example, a study by Dev et al. (2021) demonstrated how biased language in training data could lead to the propagation of racial and gender stereotypes in NLP models .
@@
    • Contextual Analysis: While lexical analysis identifies the presence of specific biased terms, contextual analysis delves deeper into how these terms are used within the text. This method employs advanced NLP techniques to examine the context in which potentially biased language appears, allowing for the identification of subtler forms of bias. For instance, the same word might be used neutrally in one context but carry a prejudiced meaning in another. Contextual analysis examines sentence structure, co-occurrence patterns, and the surrounding language to uncover these nuances. Studies like those by Bolukbasi et al. (2016) have shown that even when biased language is not overtly present, underlying patterns can still perpetuate stereotypes, such as gender biases in word embeddings . This approach is crucial for identifying and mitigating biases that may not be immediately apparent but can significantly impact the behavior of LLMs. Zhao et al. (2019) quantified and analyzed gender bias exhibited in ELMo’s contextualized word vectors. contextual analysis can be implemented by methods like template-based method Kurita et al. (2019), applied thematic analysis Mackieson et al. (2019), and contextualized word embeddings analysis Basta et al. (2019).
@@
    • Sentiment Analysis: Sentiment analysis is another vital tool in detecting bias in text data. This technique assesses the sentiment—positive, negative, or neutral—associated with different demographic groups or topics within the training data. By analyzing how different groups are described, researchers can identify patterns of negative or biased portrayals that could influence the model’s outputs. For example, if certain demographic groups are consistently associated with negative emotions or connotations, the model might learn to replicate these biases in its predictions or interactions. Research by Kiritchenko and Mohammad (2018) has highlighted how sentiment analysis can be used to detect and address such biases in text data . This method provides a quantitative measure of bias, enabling targeted interventions to reduce the impact of these biases on the model’s behavior. Sentiment analysis can be implemented by lots of traditional machine learning tools Wankhade et al. (2022); Saxena et al. (2022); Hasan et al. (2019), pre-processed data framework Hasan et al. (2019) or transformer-based methods Devlin et al. (2018); Naseem et al. (2020); Abdullah and Ahmet (2022).
@@
  • Annotation Bias Analysis: Annotation Bias Analysis is a crucial component in evaluating and mitigating bias in LLMs. This process involves examining the biases introduced during the data annotation phase, where human annotators label or categorize training data. Since annotators bring their own biases and perspectives, their subjective decisions can inadvertently introduce skewed or biased annotations, which in turn affect the performance and fairness of the LLMs. To perform annotation bias analysis, researchers must review the annotation guidelines and the training process for annotators, ensuring that they are designed to minimize bias. Additionally, evaluating the consistency and fairness of annotations across different demographic groups and annotators helps identify any disparities. Tools such as inter-annotator agreement metrics Artstein (2017) and statistical analysis Paun et al. (2022) of annotated data can reveal potential biases. For example, research by Havens et al. (2022) on gender bias in annotated datasets highlights how annotation practices can reinforce stereotypes and biases, emphasizing the need for rigorous analysis and correction methods. Another way to reduce annotation bias is human-LLM collaborative approach Wang et al. (2024b). By addressing annotation bias, researchers can enhance the quality and fairness of training data, leading to more balanced and unbiased LLMs.
@@ 4.2 Model-Level Bias Evaluation Methods
  • Fairness Metrics: These metrics are pivotal in evaluating the model’s output fairness across different groups. Common metrics include:
    • Equal Opportunity: Ensures that the True Positive Rate (TPR) is similar across groups defined by sensitive attributes such as gender or race. For example, a sentiment analysis model should equally recognize positive sentiments across both male and female-associated names Liang et al. (2021); Gallegos et al. (2024).
@@
  • Interpretability Tools: Tools such as SHAP (Shapley Additive Explanations) Mosca et al. (2022) and LIME (Local Interpretable Model-agnostic Explanations) Wu et al. (2024) provide insights into how specific features influence model predictions, helping identify biases. For instance, SHAP values can reveal that gendered words heavily influence predictions, indicating potential gender bias in the model Liang et al. (2021); Lin et al. (2024). Similarly, LIME approximates complex models locally with interpretable models to highlight the impact of features on individual predictions, which is crucial for diagnosing bias Wu et al. (2024).
  • Counterfactual Fairness: This method generates counterfactual examples by altering sensitive attributes (e.g., changing names from traditionally male to female) to see if the model’s outputs remain invariant. A fair model should ideally produce the same outcomes regardless of these attribute changes Doan et al. (2024); Li et al. (2023).
Table 3: Summary of Model-Level Bias Evaluation Methods
@@
Fairness Metrics Equal Opportunity: Ensures similar True Positive Rates (TPR) across sensitive groups.
@@ 4.3 Output-Level Bias Evaluation Methods
Output-level bias evaluation methods assess how LLMs generate responses, especially with regard to fairness and neutrality across different demographic groups. These methods aim to identify whether the model’s predictions, recommendations, or generated content reflect biases that may have been learned during training, including those based on race, gender, socioeconomic status, political orientation, and more. Here, we explore five major metrics used to detect bias in LLM-generated text: Counterfactual Testing, Stereotype Detection, Sentiment and Toxicity Analysis, Acceptance and Rejection Rates, and Embedding-Based Metrics.
  • Counterfactual Testing: Counterfactual testing involves modifying input prompts by altering specific attributes, such as gender, race, or ethnicity, while keeping the rest of the context unchanged. This method evaluates whether the LLM’s output varies based on these demographic changes, allowing researchers to isolate the effect of these attributes on the model’s behavior. For instance, swapping "John" with "Maria" in a sentence can reveal if the model responds differently due to the gender of the subject.
    Recent work Wang (2024) emphasizes the importance of robust causal reasoning in LLMs to enhance fairness, arguing that a strong understanding of causal relationships can mitigate biases and reduce hallucinations. Another study Banerjee et al. (2024) introduces a dynamic framework that compares outputs across different demographic groups in the same context, improving the fairness of generated text without requiring costly model retraining. Additionally, new tools Cheng et al. (2024) for generating and analyzing counterfactuals allow users to explore LLM behavior interactively, ensuring the counterfactuals are both meaningful and grammatically accurate.
@@
    One approach Wu et al. (2024) introduced a unified dataset combining multiple stereotype detection datasets. Researchers fine-tuned LLMs on this dataset and found that multi-dimensional classifiers were more effective in identifying stereotypes. This study also highlights the use of explainable AI tools to ensure models align with human understanding. Another study Bai et al. (2024) developed a dual-framework that combines static evaluations with dynamic, real-world scenario simulations. This dynamic aspect is particularly effective in detecting subtle, context-specific biases that static tests might miss. A qualitative method Babonnaud et al. (2024) uses prompting techniques to uncover implicit stereotypes in LLM-generated text. This approach, focusing on biases like gender and ethnicity, employs the Tree of Thoughts technique to systematically reveal hidden prejudices and provide a reproducible method for stereotype detection. These methods collectively help researchers detect and mitigate biases in LLM outputs, offering a deeper understanding of how stereotypes manifest in generated text.
@@
    Llama Guard Inan et al. (2023) introduces a model-based approach to bias evaluation by using a safety risk taxonomy for classifying both prompts and responses in human-AI conversations. This model, fine-tuned on a carefully curated dataset, shows strong performance in detecting various forms of toxicity. It represents a practical application of bias evaluation methods by providing a dynamic tool for assessing and mitigating harmful content in real-world interactions. The definition-based toxicity metric Koh et al. (2024) offers a flexible solution for bias evaluation by using LLMs to measure toxicity according to predefined criteria. This method outperforms traditional metrics, enhancing the F1 score significantly and addressing the limitations of existing models that rely on dataset-specific definitions. It exemplifies how bias evaluation methods can be refined to improve the detection of nuanced toxic content. Moderation Using LLM Introspection (MULI) Hu et al. (2024) advances bias evaluation by analyzing internal model responses to detect toxic prompts. This approach leverages patterns in response token logits and refusal behaviors, providing a cost-effective way to assess biases without requiring additional training.
@@
    In one study An et al. (2024), researchers assessed how LLMs respond to job applicants based on perceived race, ethnicity, and gender by manipulating first names. They found that acceptance rates were significantly higher for masculine White names compared to masculine Hispanic names. These acceptance and rejection rates were highly prompt-sensitive, meaning that variations in how the question was framed could lead to different outcomes. This suggests that LLMs can subtly reinforce biases, depending on how inputs are structured. Another evaluation method Scherrer et al. (2024) used to study LLM behavior in moral decision-making showed that models are consistent in choosing commonsense actions in unambiguous scenarios, while in more ambiguous cases, LLMs exhibited uncertainty and greater variability. Closed-source models often displayed more consistent preferences in these ambiguous scenarios, which suggests that proprietary models might encode specific tendencies in moral decision-making. BiasBuster Echterhoff et al. (2024) provides a systematic method for uncovering and evaluating cognitive biases in LLMs, particularly in high-stakes decision-making. By using a dataset of 16,800 prompts designed to probe various types of biases (e.g., prompt-induced, sequential), this framework allowed researchers to assess how LLMs handled acceptance and rejection decisions under different bias conditions.

### v2

@@ 4 Bias Evaluation
As LLMs become increasingly integrated into various real-world applications, from healthcare decision making to legal judgments and everyday digital interactions, their potential to propagate biases has raised significant ethical and societal concerns Liang et al. (2021). These biases, deeply rooted in the data, model architecture, and even in post-processing steps, can lead to discriminatory outcomes against marginalized groups, amplifying existing inequalities Gallegos et al. (2024). Therefore, understanding and evaluating these biases are crucial steps in ensuring the fair and responsible deployment of LLMs.
@@
To avoid ambiguity, we distinguish the taxonomy of bias types from the taxonomy of bias evaluation methods. The earlier intrinsic–extrinsic distinction classifies where bias originates or manifests, whereas the present section classifies how bias is evaluated. Within this evaluation taxonomy, data-level, model-level, and output-level methods refer to the primary technical levels at which bias is assessed. In contrast, human-involved and domain-specific evaluations are best understood as cross-cutting evaluation modes: they can be applied to data-, model-, or output-level analyses rather than forming mutually exclusive categories of the same type.
@@ 4.1 Data-Level Bias Evaluation Methods
Data-level bias evaluation methods focus on identifying and quantifying biases inherent in the training data of LLMs. These biases, rooted in the data, can significantly impact model outputs and exacerbate societal stereotypes. Figure 2 illustrates the taxonomy for data-level bias evaluation that we discuss in this section.
Figure 2: Taxonomy of data-level bias evaluation methods.
@@
    • Representation across demographics: Research has consistently demonstrated that the representation of demographics in training data can significantly influence the outcomes of LLMs. For instance, LLMs trained on datasets with disproportionate representation of certain demographic groups are prone to exhibit biases that favor those groups Simmons and Hare (2023); Wang et al. (2024); Gorti et al. (2024). For example, an analysis of political biases in LLMs revealed that most conversational LLMs exhibit left-of-center political preferences when probed with politically charged questions or statements Rozado (2024). The study demonstrated that LLMs can be steered towards specific political orientations through supervised fine-tuning (SFT) with modest amounts of politically aligned data.
      To assess representation, statistical tools can be employed to measure the distribution of demographic attributes within the dataset. Key metrics might include the relative frequency of each demographic group or the variance in representation across different categories. Visual tools such as histograms, bar charts, or demographic distribution tables can be useful for identifying disparities, providing a clear picture of which groups are over-represented or under-represented in the data.
@@
    • Data source analysis: The quality and bias of a model are heavily influenced by the sources of the data used for training. Biases inherent in the data sources can significantly impact the model’s behavior. For instance, if a model is trained primarily on data from Western countries, it may perform poorly when applied in non-Western contexts.
@@
      Evaluating the bias of data sources for LLM training involves analyzing the origins, diversity, and quality of the data, as well as its impact on model performance. This includes cataloging the data sources to ensure they cover a wide range of geographical, cultural, and linguistic contexts, and examining the credibility and inherent biases of each source. For instance, research Shen et al. (2023) on the SlimPajama dataset, which includes a rigorously deduplicated combination of web text, Wikipedia, GitHub, and books, reveals how different data combinations affect LLM performance. The SlimPajama-DC study highlights two key aspects: the impact of global versus local deduplication on model performance and the importance of data diversity post-deduplication. The findings indicate that models trained with highly deduplicated, diverse datasets outperform those trained on less-refined data, underscoring the significance of comprehensive and balanced data sources. However, duplicate removal is not always neutral with respect to representation. If content from minority, low-resource, or less frequently documented communities is already sparse, aggressive deduplication may disproportionately remove repeated expressions of those voices, thereby reducing their visibility in the corpus and unintentionally increasing representativeness bias. Bias detection tools and adjustments in data weighting can further help in managing these biases, supported by transparent documentation and regular reviews to ensure fairness and accuracy in the model outputs.
@@
  • Stereotype and bias detection in text data: Stereotype and bias detection involves analyzing the content of the training data to identify and quantify biases related to stereotypes, offensive language, or prejudiced statements. This can be achieved through the following.
@@
    • Lexical analysis: This method focuses on identifying specific words or phrases within the training data that are associated with bias or stereotypes. Lexical analysis relies on predefined lexicons, which are curated lists of terms known to carry biased or stereotypical connotations. Tools like Hatebase, an extensive repository of hate speech terms, provide a valuable resource for identifying harmful language in text data. Hate speech can also be measured by a large-scale study Silva et al. (2016). BiCapsHate Kamal et al. (2023) is a deep learning model used to detect hate speech in online social media posts. Similarly, the Dictionary of Affect in Language (DAL) Whissell (1989) categorizes words based on their emotional connotations, offering insights into how language can perpetuate stereotypes. By scanning the training data for occurrences of these terms, researchers can quantify the prevalence of biased language and identify specific areas where the data may reinforce harmful stereotypes. For example, a study by Dev et al. (2021) demonstrated how biased language in training data could lead to the propagation of racial and gender stereotypes in NLP models.
@@
    • Contextual analysis: While lexical analysis identifies the presence of specific biased terms, contextual analysis delves deeper into how these terms are used within the text. This method employs advanced NLP techniques to examine the context in which potentially biased language appears, allowing for the identification of subtler forms of bias. For instance, the same word might be used neutrally in one context but carry a prejudiced meaning in another. Contextual analysis examines sentence structure, co-occurrence patterns, and the surrounding language to uncover these nuances. A similar study by Bolukbasi et al. (2016) has shown that even when biased language is not overtly present, underlying patterns can still perpetuate stereotypes, such as gender biases, in word embeddings. This approach is crucial for identifying and mitigating biases that may not be immediately apparent but can significantly impact the behavior of LLMs. Zhao et al. Zhao et al. (2019) quantified and analyzed the gender bias exhibited in ELMo’s contextualized word vectors. Contextual analysis can be implemented by methods like template-based method Kurita et al. (2019), applied thematic analysis Mackieson et al. (2019), and contextualized word embedding analysis Basta et al. (2019).
@@
    • Sentiment analysis: Sentiment analysis is another vital tool in detecting bias in text data. This technique assesses the sentiment—positive, negative, or neutral—associated with different demographic groups or topics within the training data. By analyzing how different groups are described, researchers can identify patterns of negative or biased portrayals that could influence the model’s outputs. For example, if certain demographic groups are consistently associated with negative emotions or connotations, the model might learn to replicate these biases in its predictions or interactions. Research by Kiritchenko and Mohammad (2018) has highlighted how sentiment analysis can be used to detect and address such biases in text data. This method provides a quantitative measure of bias, enabling targeted interventions to reduce the impact of these biases on the model’s behavior. Sentiment analysis can be implemented using various approaches, including traditional machine learning methods Wankhade et al. (2022); Saxena et al. (2022); Hasan et al. (2019), frameworks based on preprocessed data Hasan et al. (2019), and transformer-based methods Devlin et al. (2018); Naseem et al. (2020); Abdullah and Ahmet (2022).
@@
  • Annotation bias analysis: Annotation bias analysis is a crucial component in evaluating and mitigating bias in LLMs. This process involves examining the biases introduced during the data annotation phase, where human annotators label or categorize training data. Since annotators bring their own biases and perspectives, their subjective decisions can inadvertently introduce skewed or biased annotations, which in turn affect the performance and fairness of the LLMs. To perform annotation bias analysis, researchers must review the annotation guidelines and the training process for annotators, ensuring that they are designed to minimize bias. Additionally, evaluating the consistency and fairness of annotations across different demographic groups and annotators helps identify any disparities. Tools such as inter-annotator agreement metrics Artstein (2017) and statistical analysis Paun et al. (2022) of annotated data can reveal potential biases. For example, research by Havens et al. (2022) on gender bias in annotated datasets highlights how annotation practices can reinforce stereotypes and biases, emphasizing the need for rigorous analysis and correction methods. Another way to reduce annotation bias is a human–LLM collaborative approach Wang et al. (2024). By addressing annotation bias, researchers can enhance the quality and fairness of training data, leading to more balanced and unbiased LLMs.
@@ 4.2 Model-Level Bias Evaluation Methods
  • Fairness metrics: These metrics are pivotal in evaluating the model’s output fairness across different groups. Common metrics include the following.
    • Equal opportunity: Ensures that the true positive rate (TPR) is similar across groups defined by sensitive attributes such as gender or race. For example, a sentiment analysis model should equally recognize positive sentiments across both male- and female-associated names Liang et al. (2021); Gallegos et al. (2024).
@@
  • Interpretability tools: Tools such as SHAP (Shapley Additive Explanations) Mosca et al. (2022) and LIME (Local Interpretable Model-agnostic Explanations) Wu et al. (2024) provide insights into how specific features influence model predictions, helping identify biases. For instance, SHAP values can reveal that gendered words heavily influence predictions, indicating potential gender bias in the model Liang et al. (2021); Lin et al. (2024). Similarly, LIME approximates complex models locally with interpretable models to highlight the impact of features on individual predictions, which is crucial for diagnosing bias Wu et al. (2024). However, these tools should be interpreted with caution in LLMs with billions of parameters. In such massive models, feature attributions are often approximate and local, and may be unstable across prompt wording, tokenization, sampling settings, or semantically similar inputs. They may also fail to capture complex high-order interactions and therefore should not be treated as definitive causal explanations of bias. For this reason, SHAP- and LIME-based analyses are best used as diagnostic signals that should be corroborated with counterfactual, robustness, and output-level evaluations.
  • Counterfactual fairness: This method generates counterfactual examples by altering sensitive attributes (e.g., changing names from traditionally male to female) to see if the model’s outputs remain invariant. A fair model should ideally produce the same outcomes regardless of these attribute changes Doan et al. (2024); Li et al. (2023). Here, counterfactual fairness is used in the model-level sense: it examines whether the model’s decision rule or prediction remains invariant when only a sensitive attribute is changed, rather than comparing differences in the wording or tone of free-form generated responses.
Table LABEL:tab:model_bias_evaluation_clean summarizes the three categories of model-level bias evaluation methods discussed above.
Table 3: Summary of model-level bias evaluation methods.
@@
Fairness metrics Equal opportunity: Ensures similar true positive rates (TPRs) across sensitive groups.
@@ 4.3 Output-Level Bias Evaluation Methods
Output-level bias evaluation methods assess how LLMs generate responses, especially with regard to fairness and neutrality across different demographic groups. These methods aim to identify whether the model’s predictions, recommendations, or generated content reflect biases that may have been learned during training, including those based on race, gender, socioeconomic status, political orientation, and more. Here, we explore four major metrics used to detect bias in LLM-generated text: counterfactual testing, stereotype detection, sentiment and toxicity analysis, and acceptance and rejection rates (see Figure 3). Unlike model-level counterfactual fairness, output-level counterfactual testing compares the generated outputs themselves under minimally modified prompts, focusing on whether the content, framing, sentiment, or recommendation changes across demographic substitutions.
Figure 3: Taxonomy of output-level bias evaluation methods.
  • Counterfactual testing: Counterfactual testing involves modifying input prompts by altering specific attributes, such as gender, race, or ethnicity, while keeping the rest of the context unchanged. This method evaluates whether the LLM’s output varies based on these demographic changes, allowing researchers to isolate the effect of these attributes on the model’s behavior. For instance, swapping “John” with “Maria” in a sentence can reveal if the model responds differently due to the gender of the subject.
    Recent work Wang (2024) emphasizes the importance of robust causal reasoning in LLMs to enhance fairness, arguing that a strong understanding of causal relationships can mitigate biases and reduce hallucinations. Beyond fairness, this connection is also important for understanding hallucinations. When a prompt is ambiguous or underspecified, biased social associations learned from the training corpus may act as anchors, leading the model to generate stereotype-consistent details even when those details are not supported by the input. For example, the model may infer occupations, traits, intentions, or background facts from demographic cues because such associations are statistically frequent in the data, not because they are causally justified in the given context. In this sense, some hallucinations can be viewed not merely as factual errors, but as bias-driven distortions in which stereotype-consistent priors fill evidential gaps. Stronger causal reasoning may reduce this risk by helping the model distinguish socially correlated attributes from causally relevant information. Another study Banerjee et al. (2024) introduces a dynamic framework that compares outputs across different demographic groups in the same context, improving the fairness of generated text without requiring costly model retraining. Additionally, new tools Cheng et al. (2024) for generating and analyzing counterfactuals allow users to explore LLM behavior interactively, ensuring the counterfactuals are both meaningful and grammatically accurate.
@@
    One approach Wu et al. (2024) introduced a unified dataset combining multiple stereotype detection datasets. Researchers fine-tuned LLMs on this dataset and found that multi-dimensional classifiers were more effective in identifying stereotypes. This study also highlighted the use of explainable AI tools to ensure models align with human understanding. Another study Bai et al. (2024) developed a dual framework that combines static evaluations with dynamic, real-world scenario simulations. This dynamic aspect is particularly effective in detecting subtle, context-specific biases that static tests might miss. A qualitative method presented in  Babonnaud et al. (2024) uses prompting techniques to uncover implicit stereotypes in LLM-generated text. This approach, focusing on biases like gender and ethnicity, employs the Tree of Thoughts technique to systematically reveal hidden prejudices and provide a reproducible method for stereotype detection. These methods collectively help researchers detect and mitigate biases in LLM outputs, offering a deeper understanding of how stereotypes manifest in generated text.
@@
    Llama Guard Inan et al. (2023) introduces a model-based approach to bias evaluation by using a safety risk taxonomy for classifying both prompts and responses in human–AI conversations. This model, fine-tuned on a carefully curated dataset, shows strong performance in detecting various forms of toxicity. It represents a practical application of bias evaluation methods by providing a dynamic tool for assessing and mitigating harmful content in real-world interactions. The definition-based toxicity metric Koh et al. (2024) offers a flexible solution for bias evaluation by using LLMs to measure toxicity according to predefined criteria. This method outperforms traditional metrics, enhancing the F1 score significantly and addressing the limitations of existing models that rely on dataset-specific definitions. It exemplifies how bias evaluation methods can be refined to improve the detection of nuanced toxic content. Moderation Using LLM Introspection (MULI) Hu et al. (2024) advances bias evaluation by analyzing internal model responses to detect toxic prompts. This approach leverages patterns in response token logits and refusal behaviors, providing a cost-effective way to assess biases without requiring additional training.
@@
    In one study An et al. (2024), researchers assessed how LLMs respond to job applicants based on perceived race, ethnicity, and gender by manipulating first names. They found that acceptance rates were significantly higher for masculine White names compared to masculine Hispanic names. These acceptance and rejection rates were highly prompt-sensitive, meaning that variations in how the question was framed could lead to different outcomes. This suggests that LLMs can subtly reinforce biases depending on how inputs are structured. Another evaluation method Scherrer et al. (2024) used to study LLM behavior in moral decision making showed that models are consistent in choosing commonsense actions in unambiguous scenarios, while in more ambiguous cases, LLMs exhibited uncertainty and greater variability. Closed-source models often displayed more consistent preferences in these ambiguous scenarios, which suggests that proprietary models might encode specific tendencies in moral decision making. BiasBuster Echterhoff et al. (2024) provides a systematic method for uncovering and evaluating cognitive biases in LLMs, particularly in high-stakes decision making. By using a dataset of 16,800 prompts designed to probe various types of biases (e.g., prompt-induced, sequential), this framework allowed researchers to assess how LLMs handled acceptance and rejection decisions under different bias conditions.